# Quantum optical neural networks with programmable nonlinearities

### v1

@@ Appendix B Optimization algorithm
We use gradient-free optimization algorithms that optimize an objective function without explicitly defined derivatives. We utilized Controlled Random Search (CRS) [45] as the global optimizer and Bound Optimization BY Quadratic Approximation (BOBYQA) [46] for local optimization as implemented in NLopt library [47]. CRS algorithm starts with a random population of points with size $N=10\times(P+1)$, and evolves these points by specified rules. It was chosen because of its ability to work efficiently in large-dimensional parameter spaces, which is difficult for many optimization algorithms. The size of the parameter space plays an important role in training a deep neural network. For the same reason, BOBYQA is preferred. This algorithm performs bound-constrained optimization using an iteratively constructed quadratic approximation for the objective function using $m$ iterpolation points, the value $m=2P+1$ being typical. Moreover, the combination of these two algorithms demonstrated the best convergence among other methods from the NLopt library that we have tested.

### v2

@@ Appendix B Optimization algorithm
We use gradient-free optimization algorithms that optimize an objective function without explicitly defined derivatives. We utilized Controlled Random Search (CRS) [54] as the global optimizer and Bound Optimization BY Quadratic Approximation (BOBYQA) [55] for local optimization as implemented in NLopt library [56]. CRS algorithm starts with a random population of points with size $N=10\times(P+1)$, and evolves these points by specified rules. It was chosen because of its ability to work efficiently in large-dimensional parameter spaces, which is difficult for many optimization algorithms. The size of the parameter space plays an important role in training a deep neural network. For the same reason, BOBYQA is preferred. This algorithm performs bound-constrained optimization using an iteratively constructed quadratic approximation for the objective function using $m$ interpolation points, the value $m=2P+1$ being typical. Moreover, the combination of these two algorithms demonstrated the best convergence among other methods from the NLopt library that we have tested.
It is necessary to run the optimization algorithm multiple times to achieve a minimum value of the cost function. To illustrate this, Fig. 14 shows the dependence of infidelity for the QONN with adjustable nonlinearities that prepares the $3$-qubit state $|G_{N}(\alpha)\rangle$ for $\alpha=\pi/4$. At each value depth $D_{\text{NL}}$ the optimization algorithm was ran $50$ times, each time initialized by random values of the QONN parameters. The points in the figure correspond to the infidelities obtained. The best obtained values (connected points) corresponds to one dependence from Fig. 4b.
Figure 14: Example of infidelity values obtained in multiple runs of optimizatin algorithm for the QONN generating $3$-qubit state $|G_{N}(\alpha)\rangle$ at $\alpha=\pi/4$ and $\varphi_{b}=\pi/4$. The best values (marked by connected points) to one curve depicted in Fig. 4b.
Appendix C Multi-photon dynamics inside the QONNs
Figure 15: (left column) Example of distributions of absolute values of the probability amplitudes of the quantum states entering the nonlinear layers of the QONN with adjustable nonlinearities and (right column) corresponding nonlinearity strengths within the quantum circuit. The QONN has been trained to perform $3$-qubit VQE algorithm with Heisenberg Hamiltonian on the $2D$ triangular lattice model. The QONN has $D_{\text{NL}}=3$ programmable nonlinear layers.
Figure 16: Example of distributions of absolute values of the probability amplitudes of the quantum states entering the nonlinear layers of the QONN with programmable linear optics. The QONN has been trained to perform $3$-qubit VQE algorithm with Heisenberg Hamiltonian on the $2D$ triangular lattice model. The QONN has $D_{\text{NL}}=1$ nonlinear layer with static nonlinearities $\chi=\pi$.
[…]
Fig. 15 and fig. 16 show an example of the probability amplitude distribution of the state before the nonlinear layers within neural network with programmable nonlinearities and linear optical network, respectively, for the VQE of the 3-qubit Heisenberg Hamiltonian $H$ on the 2D triangular lattice model.